OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (e.mccormick):

Ah, it is a system, so you just try points and see if they are valid or not for the system. Sometimes this is eaiest if you grap the system and look for overlap.

OpenStudy (anonymous):

(-5, -2) @e.mccormick

OpenStudy (e.mccormick):

Id "y less than or greater to x - 5" supposed to be "y less than or equal to x - 5" ?

OpenStudy (anonymous):

?

OpenStudy (e.mccormick):

Well, look at your text in the question. The first part is a bit impossible.

OpenStudy (e.mccormick):

\[y < > x-5 \\ y\ge -x-4 \] is what you wrote.

OpenStudy (anonymous):

can you tell me the answer I don't know it

OpenStudy (e.mccormick):

I just need you to clarify the question. The first line you wrote is impossible. I THINK you meant \(y\le x-5\).

OpenStudy (anonymous):

yeah that's what I meant

OpenStudy (e.mccormick):

K. Then does this help? https://www.desmos.com/calculator/w8enpxgeav

OpenStudy (e.mccormick):

The shaded regions show the graphs of each equation. When you look at an inequality like this in x and y, you want to know where the y values will be. So you find the line y=x-5, then vecause this is \(\le\) you mark out below that (less than that). Then you graph the line y=-x-4, and because this is really \(\ge\) you shade out above that (greater than). Where you have both, that is a valid solution.

OpenStudy (anonymous):

(5, -2)

OpenStudy (e.mccormick):

Yes!

OpenStudy (e.mccormick):

I hope you can see why so you can get this type of thing on your own. It is all about the graphs and understanding them.